• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name lkaufman

    Log entry time 10:05:57 on September21,2005

    Entry number 155487

    Followups:

    keyword=Cavity monitors stopped working - working again!

    The cavity monitors stopped working during run 5220.  It seems that some of the 
    cavities started working again during 5231.  The downstream bcm and x cavity 
    did not come back.  When I logged in to check the status of the cavities, the 
    epics screens were white, so I rebooted the crate.  The cavities are working 
    again.
    
    Fig. 1; Upstream cavity
    Fig. 2: Downstream cavity
    


    FIGURE 1

    FIGURE 2